OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a 12 -inch diameter pizza and an 18 -inch diameter pizza. He has set the price of an 12 -inch cheese pizza at $8.00, based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an 18 -inch cheese pizza? Show and explain your solution.

OpenStudy (anonymous):

12/18=$8.00/x So x = $8.00(18/12) So x = $12.00